# Strategic Learning and Trading in Broker-Mediated Markets

### v1

@@ 4.3   Hamilton–Jacobi–Bellman equation
$\eta^{*}=\frac{z_{1}^{I}(t)\,\hat{\alpha}+z_{2}^{I}(t)\,\nu+2\,g_{2}^{I}(t)\,q^{I}}{2\,\mathfrak{b}}\,.$ (4.43)
Next, define
and employ the ansatz
$H=x^{B}+h(t,s,q^{B},\hat{\alpha},\xi,q^{I})$
to obtain the following PDE in $h$
$\begin{split}0&=\partial_{t}h+\partial_{s}h\,\hat{\alpha}+\xi\left(s+\mathfrak{c}\,\xi\right)-\partial_{q^{B}}h\,\xi+\partial_{\hat{\alpha}}h\left(-\kappa^{\alpha}\hat{\alpha}\right)+\partial_{\xi}h\left(-\kappa^{U}\,\xi\right)\\ &\phantom{{}={}}+\frac{1}{2}\partial_{ss}h\left(\sigma^{S}\right)^{2}+\frac{1}{2}\partial_{\hat{\alpha}\hat{\alpha}}h\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\sigma^{\alpha}}{\sigma^{S}}\right)^{2}+\frac{1}{2}\partial_{\xi,\xi}h\left(\sigma^{U}\right)^{2}\\ &\phantom{{}={}}+\partial_{s\hat{\alpha}}h\left[\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}\right]-(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{t})\,\left(q^{B}\right)^{2}\\ &\phantom{{}={}}+\left[\hat{\alpha}\,f_{1}+f_{3}\,q^{I}\right]\partial_{q^{I}}h-\left[\hat{\alpha}\,f_{1}+f_{3}\,q^{I}\right]\,\partial_{q^{B}}h\\ &\phantom{{}={}}+\Bigg{[}\hat{\alpha}^{2}\,f_{1}^{2}\,\mathfrak{b}+2\,\hat{\alpha}\,f_{1}\,f_{3}\,\mathfrak{b}\,q^{I}+f_{3}^{2}\,\mathfrak{b}\,\left(q^{I}\right)^{2}+\hat{\alpha}\,f_{1}\,s+f_{3}\,q^{I}\,s\Bigg{]}\\ &\phantom{{}={}}+\sup_{\nu}\Bigg{\{}-\left(\nu\right)^{2}\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)+\nu\bigg{[}f_{2}\,\partial_{q^{I}}h+\left(1-f_{2}\right)\,\partial_{q^{B}}h\\ &\phantom{{}={}}\qquad+\left(2\,\hat{\alpha}\,f_{1}\,f_{2}\,\mathfrak{b}+2\,f_{2}\,f_{3}\,\mathfrak{b}\,q^{I}-s+f_{2}\,s\right)\,+\mathfrak{p}\,\partial_{s}h\bigg{]}\Bigg{\}}\,,\end{split}$ (4.44)
with terminal condition
The supremum in (4.44) is attained at
$\nu^{*}=\frac{f_{2}\,\partial_{q^{I}}h+\left(1-f_{2}\right)\,\partial_{q^{B}}h+\left(2\,\hat{\alpha}\,f_{1}\,f_{2}\,\mathfrak{b}+2\,f_{2}\,f_{3}\,\mathfrak{b}\,q^{I}-s+f_{2}\,s\right)+\mathfrak{p}\,\partial_{s}h}{2\,\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)}\,,$ (4.45)
$\begin{split}0&=\partial_{t}h+\partial_{s}h\,\hat{\alpha}+\xi\,\left(s+\mathfrak{c}\,\xi\right)-\partial_{q^{B}}h\,\xi+\partial_{\hat{\alpha}}h\left[-\kappa^{\alpha}\,\hat{\alpha}\right]+\partial_{\xi}h\left[-\kappa^{U}\,\xi\right]\\ &\phantom{{}={}}+\frac{1}{2}\partial_{ss}h\left(\sigma^{S}\right)^{2}+\frac{1}{2}\partial_{\hat{\alpha}\hat{\alpha}}h\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}}{\sigma^{S}}\right)^{2}+\frac{1}{2}\partial_{\xi,\xi}h\left(\sigma^{U}\right)^{2}\\ &\phantom{{}={}}+\partial_{s\hat{\alpha}}h\left[\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\sigma^{\alpha}\right]-(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{t})\,q^{2}\\ &\phantom{{}={}}+\left[\hat{\alpha}\,f_{1}+f_{3}\,q^{I}\right]\partial_{q^{I}}h-\left[\hat{\alpha}\,f_{1}+f_{3}\,q^{I}\right]\,\partial_{q^{B}}h\\ &\phantom{{}={}}+\Bigg{[}\hat{\alpha}^{2}\,f_{1}^{2}\,\mathfrak{b}+2\,\hat{\alpha}\,f_{1}\,f_{3}\,\mathfrak{b}\,q^{I}+f_{3}^{2}\,\mathfrak{b}\,\left(q^{I}\right)^{2}+\hat{\alpha}\,f_{1}\,s+f_{3}\,q^{I}\,s\Bigg{]}\\ &\phantom{{}={}}+\frac{\left[f_{2}\,\partial_{q^{I}}h+\left(1-f_{2}\right)\,\partial_{q^{B}}h+\left(2\,\hat{\alpha}\,f_{1}\,f_{2}\,\mathfrak{b}+2\,f_{2}\,f_{3}\,\mathfrak{b}\,q^{I}-s+f_{2}\,s\right)+\mathfrak{p}\,\partial_{s}h\right]^{2}}{4\,\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)}\,.\end{split}$ (4.46)
The HJB equation in matrix form
To study the problem further, we show that the PDE (4.46) is a system of matrix differential equations involving a matrix Riccati ODE. Employ the ansatz
$h(t,s,\tilde{\mathfrak{y}})=q^{B}\,s+G_{0}+2\,G_{1}\,\tilde{\mathfrak{y}}^{\top}+\tilde{\mathfrak{y}}\,G_{2}\,\tilde{\mathfrak{y}}^{\top}$ (4.47)
where $\tilde{\mathfrak{y}}=\begin{pmatrix}q^{B}&\hat{\alpha}&\xi&q^{I}\end{pmatrix}$, $G_{0}:[0,T]\to\mathbb{R}$, $G_{1}:[0,T]\to\mathbb{R}^{1\times 4}$, and $G_{2}:[0,T]\to\mathbb{R}_{\text{symmetric}}^{4\times 4}$ are functions of time, and write the HJB equation (4.46) as
00 0 $=$ $\Bigg{[}G_{0}^{\prime}+4\,G_{1}\,P_{8}^{\top}\,P_{8}\,G_{1}^{\top}+\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}}{\sigma^{S}}\right)^{2}\,e_{2}\,G_{2}\,e_{2}^{\top}+(\sigma^{U})^{2}\,e_{3}\,G_{2}\,e_{3}^{\top}\Bigg{]}$ (4.48)
[…]
where
$P_{2}$ $=\left(\begin{array}[]{ccccc}0&0&0&0\\ -{f_{1}}&-{\kappa^{\alpha}}&0&{f_{1}}\\ -1&0&-{\kappa^{U}}&0\\ -{f_{3}}&0&0&{f_{3}}\\ \end{array}\right)\,,$ (4.55)
$P_{5}$ $=\left(\begin{array}[]{ccccc}-{\mathbb{V}^{B}}{\rho_{1}^{B}}-{\rho_{0}^{B}}&1/2&0&0\\ 1/2&{f_{1}}^{2}{\mathfrak{b}}&0&{f_{1}}{f_{3}}{\mathfrak{b}}\\ 0&0&{\mathfrak{c}}&0&\\ 0&{f_{1}}{f_{3}}{\mathfrak{b}}&0&{f_{3}}^{2}{\mathfrak{b}}&\\ \end{array}\right)\,,$ (4.60)
$P_{7}$ $=\left(\begin{array}[]{ccccc}\frac{\mathfrak{p}}{2\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}&\frac{{f_{1}}{f_{2}}{\mathfrak{b}}}{\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}&0&\frac{{f_{2}}{f_{3}}{\mathfrak{b}}}{\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}\\ \end{array}\right)\,,$ (4.62)
$P_{8}$ $=\left(\begin{array}[]{ccccc}\frac{1-{f_{2}}}{2\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}&0&0&\frac{{f_{2}}}{2\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}\\ \end{array}\right)\,,$ (4.64)
and $\{e^{i}\}$ is the canonical basis of $\mathbb{R}^{4}$.
Recall that $\tilde{\mathfrak{y}}\,M\,\tilde{\mathfrak{y}}^{\top}=\tilde{\mathfrak{y}}\,M^{\top}\,\tilde{\mathfrak{y}}^{\top}$ for any four-by-four matrix $M$ and define
$P_{9}=2\,P_{8}^{\top}\,P_{7}+P_{2}^{\top}\,,$ (4.65)
to obtain the matrix differential equations
00 0 $=$ $\,G_{0}^{\prime}+4\,G_{1}\,P_{8}^{\top}\,P_{8}\,G_{1}^{\top}+\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}}{\sigma^{S}}\right)^{2}\,e_{2}\,G_{2}\,e_{2}^{\top}+(\sigma^{U})^{2}\,e_{3}\,G_{2}\,e_{3}^{\top}\,,$ (4.66)
00 0 $=$ $\,G_{1}^{\prime}+G_{1}\,P_{2}^{\top}+2\,G_{1}\,P_{8}^{\top}\,P_{7}+4\,G_{1}\,P_{8}^{\top}\,P_{8}\,G_{2}\,,$ (4.67)
00 0 $=$ $\,G_{2}^{\prime}+P_{7}^{\top}\,P_{7}+4\,G_{2}^{\top}\,P_{8}^{\top}\,P_{8}\,G_{2}+G_{2}^{\top}\,P_{9}+P_{9}^{\top}\,G_{2}+P_{5}\,,$ (4.68)
with terminal conditions
$G_{0}=0\,,\quad G_{1}=0\,,$
and
$G_{2}(T)=\begin{pmatrix}-(\beta_{0}^{B}+\beta_{1}^{B}\,\mathbb{V}^{B}_{T})&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}\,.$ (4.69)
The next lemma shows existence and uniqueness of a solution to the differential equation in $G_{2}$. Here, and in the remainder of the paper we assume that the values of the model parameters are such that $01+\mathfrak{b}\,f_{3}(t)>$ 0 for all $t\in[0,T]$. Note that $\mathfrak{b}\,f_{3}/2=g_{2}^{I}$, thus, the inequality can be written as $g_{2}^{I}>-1/2$. Given $g_{2}^{I}$ satisfies the ODE in (3.12), we see that if the value of the temporary impact and terminal inventory risk parameters for the informed trader ($\mathfrak{b}$, $\beta_{0}^{I}$, and $\beta_{1}^{I}$) are small, this condition is met. In the numerical examples below the values of the model parameters are such that this condition is satisfied.
Proposition 2.
There exists $\delta>0$ such that for all $\mathfrak{p}\in[0,\delta)$, the matrix Riccati equation (4.68) admits a unique solution $G_{2}:[0,T]\to\mathbb{R}^{4\times 4}$ with terminal condition (4.69).
Proof.
The matrix Riccati equation in (4.68) can be simplified because the components of $P_{8}^{\top}\,P_{8}$ are zero except in the corners, and the first and last columns of $P_{9}$ take the form $\begin{pmatrix}\cdots&0&0&\cdots\end{pmatrix}^{\top}$. The simplified differential equation satisfied by $\tilde{G}_{2}$ is
$\begin{split}0=\tilde{G}_{2}^{\prime}+4\,\tilde{G}_{2}^{\top}\,U\,\tilde{G}_{2}+\tilde{G}_{2}^{\top}\,V+V^{\top}\,\tilde{G}_{2}+B\,,\quad\tilde{G}_{2}(T)=\begin{pmatrix}-(\beta_{0}^{B}+\beta_{1}^{B}\,\mathbb{V}^{B}_{T})&0\\ 0&0\end{pmatrix}\,,\end{split}$ (4.70)
with
$\begin{split}\tilde{G}_{2}&=\begin{pmatrix}(G_{2})_{11}&(G_{2})_{14}\\ (G_{2})_{14}&(G_{2})_{44}\end{pmatrix}\,,\\ U&=\frac{1}{\mathfrak{a}-f_{2}^{2}\mathfrak{b}}\,\begin{pmatrix}(1-f_{2})^{2}&f_{2}\,(1-f_{2})\\ f_{2}\,(1-f_{2})&f_{2}^{2}\end{pmatrix}=\begin{pmatrix}1-f_{2}\\ f_{2}\end{pmatrix}\,\left(\mathfrak{a}-f_{2}^{2}\mathfrak{b}\right)^{-1}\,\begin{pmatrix}1-f_{2}&f_{2}\end{pmatrix}\,,\\ V&=\frac{1}{\mathfrak{a}-f_{2}^{2}\mathfrak{b}}\,\begin{pmatrix}\frac{\mathfrak{p}\,(1-f_{2})}{2}&-f_{3}\,(\mathfrak{a}-f_{2}\,\mathfrak{b})\\ \frac{\mathfrak{p}\,f_{2}}{2}&\mathfrak{a}\,f_{3}\end{pmatrix}\,,\\ B&=\frac{1}{\mathfrak{a}-f_{2}^{2}\mathfrak{b}}\,\begin{pmatrix}\frac{\mathfrak{p}^{2}}{4}-\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)\left(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{t}\right)&\frac{\mathfrak{p}\,f_{2}\,f_{3}\,\mathfrak{b}}{2}\\ \frac{\mathfrak{p}\,f_{2}\,f_{3}\,\mathfrak{b}}{2}&f_{3}^{2}\,\mathfrak{a}\,\mathfrak{b}\end{pmatrix}\,.\end{split}$ (4.71)
Recall that $z_{2}^{I}$ satisfies (3.27), so $f_{2}(t)=\mathfrak{p}\,\int_{t}^{T}\exp\left\{\int_{t}^{s}\frac{f_{3}(u)}{2}-\theta^{B}\,\mathrm{d}u\right\}\,\mathrm{d}s$, $f_{2}=0$ whenever $\mathfrak{p}=0$, and $f_{2}$ is continuous in $\mathfrak{p}$. Next, use Theorem 2.3 in Freiling et al., (2000) by setting
and consider the following matrices:
$B_{11}$ $=V\,,\quad B_{12}=U\,,\quad B_{21}=-B^{\top}\,,\quad B_{22}=-V^{\top}\,,$ (4.72)
$E$ $=C+D\,\tilde{G}_{2}(T)+\left(D\,\tilde{G}_{2}(T)\right)^{\top}\,,$ (4.73)
$L(t)$ $=\begin{pmatrix}C\,B_{11}+D\,B_{21}&C\,B_{12}+B_{11}^{\top}\,D+D\,B_{22}\\ {0}&B_{12}^{\top}\,D\end{pmatrix}\,$ (4.74)
$=\begin{pmatrix}C\,V+B^{\top}&C\,U\\ {0}&-U^{\top}\end{pmatrix}\,.$ (4.75)
It follows that
$E=\begin{pmatrix}2\,(\beta_{0}^{B}+\beta_{1}^{B}\,\mathbb{V}^{B}_{t})&0\\ 0&1\end{pmatrix}>0\,,\quad\text{and}\quad L+L^{\top}=\begin{pmatrix}C\,V+V^{\top}\,C+2\,B&C\,U\\ U\,C&-2\,U\end{pmatrix}\,.$ (4.76)
[…]
Lastly, observe that the differential equation for $G_{1}$ is a factor of itself and the terminal condition is $\textbf{0}\in\mathbb{R}^{1\times 4}$. Thus, the vector-valued solution is $\textbf{0}^{1\times 4}$. The ODE satisfied by $G_{0}$ is linear so it has a unique solution.
4.4   The optimal adaptive strategy
The following theorem solves the adaptive control problem of the broker.
Theorem 2.
Let $\mathfrak{p}\in[0,\delta)$ with $\delta$ as in Proposition 2. Then, the HJB equation (4.42) admits a unique solution in $C^{1,2}([0,T],\mathbb{R}^{|\mathfrak{y}|})$ with quadratic growth in $\mathfrak{y}$. Moreover, the broker’s Markovian optimal trading rate for the control problem (4.36) in feedback form is given by
where $\tilde{\mathfrak{y}}_{t}=\begin{pmatrix}Q^{B*}_{t}&\hat{\alpha}_{t}&\xi_{t}&Q^{I}_{t}\end{pmatrix}$.
Proof.
By Proposition 2 there is a unique solution to the matrix Riccati differential equation (4.68). Armed with $G_{2}$, we find the unique solution $G_{0}$ to the linear differential equation (4.66). With $G_{0}$ and $G_{2}$, we obtain the following solution to the HJB equation (4.42):
$H^{B}(t,\mathfrak{y})=x^{B}+q^{B}\,s+G_{0}(t)+\tilde{\mathfrak{y}}\,G_{2}(t)\,\tilde{\mathfrak{y}}^{\top}\,.$ (4.78)
We observe that $H^{B}$ is a quadratic polynomial in $\mathfrak{y}$ with time-dependent $C^{1}([0,T],\mathbb{R})$ coefficients; thus, it is bounded by a quadratic growth. By Theorem 3.5.2 in Pham, (2009), the broker’s strategy (4.77) is indeed her optimal Markovian control. ∎
@@ 5 Numerical results
We employ the following values of model parameters that largely follow Cartea and Sánchez-Betancourt, (2022). For the price process, the value of the permanent price impact parameter is $\mathfrak{p}=10^{-3}$, the volatility is $\sigma^{S}=1$, and the initial value of the price is $S_{0}=100$. The parameter of the signal are as follows: the mean-reversion speed is $\kappa^{\alpha}=5$, the noise of the signal $\sigma^{\alpha}=1$, and the initial value is $\alpha_{0}=0$. The correlation between the noise in the price and the signal is zero ($\rho=0$). For the uninformed trading flow, we set $\kappa^{U}=15$, $\sigma^{U}=100$, and $\xi_{0}=0$. The temporary price impacts for the broker, informed, and uninformed trader are $\mathfrak{a}=2.1\times 10^{-3}$, $\mathfrak{b}=2\times 10^{-3}$, and $\mathfrak{c}=2\times 10^{-3}$, respectively.
@@
In the informed trader’s learning equations, we set $\sigma^{B}=60$ and $\theta^{B}=10$. We assume that the initial estimate $\hat{\nu}_{0}$ is zero and we set the initial conditional variance $\mathbb{V}^{I}_{0}=0$. In the broker’s learning equations, we assume that she knows the values of the true parameters that describe the dynamics of $\alpha$, and similarly as in the informed trader’s framework, she sets $\alpha_{0}=0$ and $\mathbb{V}^{B}_{0}=0$. Both agents employ similar risk aversion parameters; we set $\beta_{0}^{I}=\beta_{0}^{B}=10^{-1}$, $\beta_{1}^{I}=\beta_{1}^{B}=10^{-3}$, $\rho_{0}^{I}=\rho_{0}^{B}=10^{-3}$, and $\rho_{1}^{I}=\rho_{1}^{B}=10^{-5}$. We run numerical simulations where we partition the time horizon $[0,T]$, with $T=1$, into $1,000$ equally-spaced intervals and we run $10,000$ simulations. [^8] The code is available at https://github.com/muhammadalifaqsha/broker_informed_noise_filtering_game.
Our parameters satisfy the conditions of Proposition 2 for existence and uniqueness of a solution. More precisely, Figure 1 shows the evolution of the first three eigenvalues of $L(t)+L(t)^{\top}$. Observe that the first three eigenvalues are negative, while the fourth eigenvalue is zero, as required in the proof of Proposition 2.
@@
Figure 2: One sample path and $90\%$ confidence bands across simulations for the following processes. The midprice $S_{t}$, the signal $\alpha_{t}$, the broker’s trading speed $\nu_{t}$, the trading speed $\eta_{t}$ of the informed trader, the trading speed $\xi_{t}$ of the uninformed trader, the cash process $X^{I}$ of the informed trader, the cash process $X^{B}$ of the broker, the inventory $Q^{I}$ of the informed trader, and the inventory $Q^{B}$ of the broker.
Figure 2 shows one sample path for (i) the midprice $S_{t}$, (ii) the signal $\alpha_{t}$, (iii) the trading speeds of the broker, the informed, and the uninformed trader, (iv) the cash processes of the informed trader and the broker, and (vi) the inventory of both the informed trader and the broker. For each process, we plot the $90\%$ confidence bands across simulations.
@@
Next, Figure 3 shows the estimates $\hat{\alpha}$ and $\hat{\nu}$. We plot the $90\%$ confidence bands across simulations. The bands around the filter are not the conditional variance of the filter, instead, they are the $90\%$ confidence bands for the filtered processes across simulations.
@@
Figure 3: One realisation of filtering carried by both informed trader and broker.
@@
Observe that the estimation accuracy of both filters is poor. This is ultimately due to the level of the permanent impact $\mathfrak{p}$ of the broker’s trading activity, and the level of the private signal $\alpha$ of the informed trader. The accuracy of our estimators are in line with practice; in financial markets, both agents face a low signal-to-noise ratio when using prices as a source of information to filter their competitor’s private information.
Lastly, Figure 4 shows the four components of the trading strategy of the broker. The main drivers are (i) inventory management (left panel), (ii) unwinding the uninformed order flow (third panel), and (iii) adjusting to the inventory level of the informed trader (fourth panel). In our simulation example, although the uninformed trading flow is profitable for the broker on average, the broker externalises part of it as seen from comparing third panel of Figure 4 and the path of $\xi$ in Figure 2. In particular, externalisation becomes more aggressive near the terminal time $T$ to reduce inventory risks. The rest of the flow is internalised.
@@
Moreover, the last panel of Figure 4 and the last panel of Figure 2 show that the broker trades against the inventory level of the informed trader, in particular, she externalises a significant portion of the informed trading flow systematically. The second panel of Figure 4 shows the speculative component of the broker’s strategy which employs the estimated signal $\hat{\alpha}$. Observe that $\hat{\alpha}$ does not contribute to the broker’s strategy to the same extent as the other three components due to the poor estimation accuracy of the filter. Finally, the first panel of Figure 4 shows that the broker’s strategy unwinds the accumulated inventory (see last panel of Figure 2).
Figure 4: The four additive components of the broker’s strategy.
5.1   Performance and comparison against market practices
We compare the performance of the broker’s strategy in (4.77) with the following benchmarks. [^9] These benchmarks are those introduced in Cartea and Sánchez-Betancourt, (2022).

### v2

@@ 4.2.2 Solving the control problem when the broker filters the trading speed
${\mathfrak{h}}_{i}(t)=\frac{h_{1,i}(t)+{\mathfrak{G}}_{1}(t)\,h_{3,i}(t)}{\mathfrak{a}},\quad i=0,\dots,5$ (4.24)
and the functions $h_{i,j}$ are given in B.7.
@@ 5 Numerical results
We employ the following values of model parameters that largely follow Cartea and Sánchez-Betancourt (2025). For the price process, the value of the permanent price impact parameter is $\mathfrak{p}=10^{-3}$, the volatility is $\sigma^{S}=1$, and the initial value of the price is $S_{0}=100$. The parameter of the signal are as follows: the mean-reversion speed is $\kappa^{\alpha}=5$, the noise of the signal $\sigma^{\alpha}=1$, and the initial value is $\alpha_{0}=0$. The correlation between the noise in the price and the signal is zero ($\rho=0$). For the uninformed trading flow, we set $\kappa^{U}=15$, $\sigma^{U}=100$, and $\xi_{0}=0$. The temporary price impacts for the broker, informed, and uninformed trader are $\mathfrak{a}=2.1\times 10^{-3}$, $\mathfrak{b}=2\times 10^{-3}$, and $\mathfrak{c}=2\times 10^{-3}$, respectively.
@@
In the informed trader’s learning equations, we set $\sigma^{B}=60$ and $\theta^{B}=10$. We assume that the initial estimate $\hat{\nu}_{0}$ is zero and we set the initial conditional variance $\mathbb{V}^{I}_{0}=0$. In the broker’s learning equations, we use the values of the true parameters that describe the dynamics of $\alpha$, and she sets $\alpha_{0}=0$ and $\mathbb{V}^{B}_{0}=0$. Both agents employ similar risk aversion parameters; we set $\beta_{0}^{I}=\beta_{0}^{B}=10^{-1}$, $\beta_{1}^{I}=\beta_{1}^{B}=10^{-3}$, $\rho_{0}^{I}=\rho_{0}^{B}=10^{-3}$, and $\rho_{1}^{I}=\rho_{1}^{B}=10^{-5}$. In 5.5, we discuss how an equilibrium of beliefs could be reached between the two players. We run numerical simulations where we partition the time horizon $[0,T]$, with $T=1$, into $1,000$ equally-spaced intervals and we run $10,000$ simulations. [^8] The code is available at https://github.com/muhammadalifaqsha/broker_informed_noise_filtering_game. Our parameters satisfy the conditions of Proposition 2 for existence and uniqueness of a solution. For more details see B.6.
Figure 1 shows, for each learning strategy, one sample path for (i) the midprice $S_{t}$, (ii) the signal $\alpha_{t}$, (iii) the trading speeds of the broker, the informed, and the uninformed trader, (iv) the cash processes of the informed trader and the broker, and (vi) the inventory of both the informed trader and the broker. For each process, we plot the $90\%$ confidence bands across simulations. As expected, for both strategies, the trajectory of the inventory processes finish near zero so agents avoid paying the terminal inventory penalties. On the other hand, when the broker employs the alternative learning strategies, her trading speeds resemble those of the informed trader. This is due to the role of the signal that plays into both players’ strategies; indeed, in the first strategy, the broker’s estimates of the signal are quite poor. We also find that if one were to use the first strategy while employing the alternative estimates of $\alpha$, the resulting strategy will resemble the second strategy.
@@
Figure 1: One sample path and $90\%$ confidence bands across simulations for the following processes. The broker’s trading speed $\nu_{t}$, the trading speed $\eta_{t}$ of the informed trader, the cash process $X^{I}$ of the informed trader, the cash process $X^{B}$ of the broker, the inventory $Q^{I}$ of the informed trader, and the inventory $Q^{B}$ of the broker. The top panel is when the broker filters from prices and the bottom panel is when she filters from the informed trader’s flow.
@@
Next, Figure 2 shows the estimates $\hat{\alpha}$ and $\hat{\nu}$ constructed by the broker and the informed trader, respectively. The top panels are for when the broker filters prices and the bottom panels are for when she filters the informed trader’s flows. We plot the $90\%$ confidence bands across simulations. The bands around the filter are not the conditional variance of the filter, instead, they are the $90\%$ confidence bands for the filtered processes across simulations.
@@
Figure 2: One realisation of filtering carried by both informed trader and broker. The top panel is when the broker filters from prices and the bottom panel is when she filters from the informed trader’s flow.
@@
For both cases, the informed trader has poor estimates of the broker’s activity in the lit market. This is ultimately due to the level of the permanent impact $\mathfrak{p}$ of the broker’s trading activity. On the other hand, the broker has poor estimates of the signal when she filters from prices, but she predicts the signal almost perfectly when she filters from the informed trader’s flow. The accuracy of our estimators are in line with practice; in financial markets, both agents face a low signal-to-noise ratio when using prices as a source of information to filter their competitor’s private information.
Lastly, Figure 3 shows the components of the trading strategy of the broker. The main drivers are (i) inventory management (left panel), (ii) unwinding the uninformed order flow (third panel), and (iii) adjusting to the inventory level of the informed trader (fourth panel). When the broker is filtering from the informed trader’s strategy, she includes an additional component $\gamma$, which is the informed trader’s flows after removing his inventory risk aversion component. As shown in the figure, the effect of $\gamma$ is negligible compared to the other four components. This shows that after correctly estimating the signal, the broker does not externalise the informed trader’s flow.
Figure 3: The additive components of the broker’s strategy. The function $h_{(\cdot)}^{B}$ refers to the coefficient of component $(\cdot)$ in the strategy; the superscript “alt” refers to the second strategy with the alternative filter. The first row is showing the four additive components when the broker filters from prices. The second and third row are showing the five additive components when the broker filters from the informed trader’s flow.
@@
Moreover, the last column of Figure 3 and the last panel of Figure 1 show that the broker trades against the inventory level of the informed trader, in particular, she externalises a significant portion of the informed trading flow systematically. The second column of Figure 3 shows the speculative component of the broker’s strategy which employs the estimated signal $\hat{\alpha}$. Observe that $\hat{\alpha}$ does not contribute to the broker’s strategy to the same extent as the other three components due to the poor estimation accuracy of the filter. On the other hand, the speculative component from exploiting $\hat{\alpha}^{\text{alt}}$ is more predominant than in the first strategy.
5.1  Performance and comparison against benchmarks
Below, we use $Q_{t}^{B,(i)}$, for $i\in\{1,2,3\}$, to denote the inventory of the broker when she follows the benchmark strategy $\nu_{t}^{(i)}$. We employ the following benchmarks.